OpenStudy (anonymous):

What is the inverse of the given relation? y = 3x + 9 A. y = 1/3x + 3 B. y = 3x - 3 C. y = 3x + 3 D. y = 1/3x - 3

OpenStudy (mertsj):

Interchange the x and y and solve the resulting equation for y

OpenStudy (anonymous):

i'm sorry i don't understand?

OpenStudy (mertsj):

Put y where x is and x where y is.

OpenStudy (mertsj):

x=3y+9

OpenStudy (mertsj):

Now solve for y

OpenStudy (anonymous):

how do i do that? subtract 9 from both sides?

OpenStudy (anonymous):

so how would i set all that up?

OpenStudy (anonymous):

First, you need to interchange x and y in the given relation which would give you: x = 3y + 9 now x - 9 = 3y or 3y = x- 9 Therefore, y = (1/3)x - (9/3) y = (1/3)x - 3 Got it? :)

OpenStudy (anonymous):

ohhh okay , thank you!! :)